OpenStudy (anonymous):

A good thinking problem. Find the number of ways in putting n different objects in m different boxes. (any number of objects in any box)

OpenStudy (experimentx):

are objects and boxes identical??

OpenStudy (anonymous):

Nope. Distinct objects. And all boxes different too.

OpenStudy (anonymous):

\( m^n \)

OpenStudy (anonymous):

@FoolForMath I thought the same. But apparently in the answer ( which is a series) some cases have been deducted.

OpenStudy (anonymous):

@siddhantsharan: what do you mean?

OpenStudy (anonymous):

As in the anser is m^n - (something)

OpenStudy (anonymous):

*Answer

OpenStudy (anonymous):

Okay then your question is not stated correctly.

OpenStudy (anonymous):

you are subtracting that means no boxes could be left empty.

OpenStudy (anonymous):

Are you sure as to subtracting being incorrect. Also there was a hint : [ Think w.r.t. inculsion exclusion principle] Annnd how did you get the second series?

OpenStudy (anonymous):

The series you are looking for is \[ \sum_{j=0}^m (-1)^{j}{m \choose j} (m-j)^n \]

OpenStudy (anonymous):

I am very sure. This is elementary mutual inclusion-exclusion. Expand that series you will get what you are looking for.

OpenStudy (anonymous):

Okay. Thanks a lot. But why that serries. How is it derieved?

OpenStudy (anonymous):

How do you derive such series?

OpenStudy (anonymous):

Annnd this subtraction (of that series) is to my question or is it when no box is empty?

OpenStudy (anonymous):

Hiatus. Lunch time :)

OpenStudy (anonymous):

im trying to understand the question. Is this the same as asking how many solutions does the equation:\[x_1+x_2+x_3+\cdots+x_m=n\]have, where each x_i must be non-negative? (think of each box as one of the variables, and the sum of the number of objects in each box must be n).

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

Wow, joemath. I love the way you think.

OpenStudy (anonymous):

If that is the case, then the answer is:\[\left(\begin{matrix}n+m-1 \\ m-1\end{matrix}\right)\]

OpenStudy (anonymous):

That is not the case :)

OpenStudy (anonymous):

How that is not the case? wait... yeah objects aren't identical.

OpenStudy (anonymous):

That's right :)